### v1

@@ Abstract
Zoned Namespace (ZNS) defines a new abstraction for host software to flexibly manage storage in flash-based SSDs as append-only zones. It also provides a Zone Append primitive to further boost the write performance of ZNS SSDs by exploiting intra-zone parallelism. However, making Zone Append effective for reliable and scalable storage, in the form of a RAID array of multiple ZNS SSDs, is non-trivial since Zone Append offloads address management to ZNS SSDs and requires hosts to dedicatedly manage RAID stripes across multiple drives. We propose ZapRAID, a high-performance log-structured RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write parallelism and lightweight stripe management. ZapRAID adopts a group-based data layout with a coarse-grained ordering across multiple groups of stripes, such that it can use small-size metadata for stripe management on a per-group basis under Zone Append. It further adopts hybrid data management to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Append and Zone Write primitives. We evaluate ZapRAID using microbenchmarks, trace-driven experiments, and real-application experiments. Our evaluation results show that ZapRAID achieves high write throughput and maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 1 Introduction
ZNS SSDs support two write primitives, namely Zone Write and Zone Append, where Zone Append provides opportunities for further performance gains by exploiting intra-zone parallelism. Specifically, a ZNS SSD tracks per-zone write pointers, such that any write to a zone must specify the same offset indicated by the write pointer. Zone Write requires the host to specify the block address when writing a block as in conventional SSDs. This provides backward compatibility with existing applications. However, in order to match the offset of a Zone Write with the on-device write pointer of a zone, the host can only issue one request to each zone at a time, thereby limiting intra-zone parallelism. In contrast, Zone Append eliminates the need of the host to specify block addresses in writes by fully offloading the address management to ZNS SSDs. In Zone Append, the host only specifies the zone to which a write is issued, and the ZNS SSD returns the address to the host upon the write completion. Thus, the host can issue multiple Zone Append commands within one zone to exploit the intra-zone parallelism for improved write performance.
Despite the performance gains from Zone Append, its offloading of address management to ZNS SSDs implies that the host not only cannot directly specify the block addresses of writes in a zone, but it also cannot control the ordering of writes in concurrent Zone Append commands. This creates new challenges to storage system designs that require host-level address management, particularly when applying Redundant Array of Independent Disks (RAID) [60] to form an array of multiple ZNS SSDs for reliable and scalable storage. Specifically, in a traditional RAID array, a RAID controller runs atop multiple drives and organizes data in stripes, each of which encodes a set of data blocks into parity blocks and distributes the data and parity blocks across drives. For efficient data repair, the RAID controller statically assigns the same block address in each drive for the blocks of the same stripe, so that the repair of any lost block can directly retrieve the available data and parity blocks of the same stripe from other drives for decoding. However, under Zone Append, it is infeasible for the RAID controller to manage block addresses in a zone; instead, the RAID controller needs to maintain dedicated address mapping information to specify the block locations for each stripe, which unavoidably incurs performance penalties to stripe management.
@@
We observe that Zone Append effectively exploits intra-zone parallelism when a workload is dominated by small writes and the writes are issued to a single zone of a ZNS SSD at a time, while Zone Write is beneficial when a workload is dominated by large writes or when the writes are concurrently issued to multiple zones of a ZNS SSD (in which Zone Write exploits better inter-zone parallelism for large writes). To this end, ZapRAID adopts hybrid data management to manage small and large writes across multiple zones in a ZNS SSD, so as to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Append and Zone Write primitives.
@@
  • We design ZapRAID, a high-performance Log-RAID system for ZNS SSDs. ZapRAID builds on two core ideas: (i) a group-based data layout under Zone Append (§3.2) and (ii) hybrid data management (§3.3) under a combination of Zone Append and Zone Write. ZapRAID further addresses reliability in terms of metadata persistence and crash consistency (§3.4).
  • We prototype ZapRAID with the Storage Performance Development Kit (SPDK) [10], such that ZapRAID can parallelize its internal operations with multiple SPDK threads.
  • We evaluate our ZapRAID prototype on real ZNS SSD devices using microbenchmarks, trace-driven experiments, and real-application experiments. When writes are issued to a single zone per ZNS SSD, ZapRAID increases the write throughput by up to 77.2% and reduces the tail latency by up to 36.4% compared with the exclusive use of Zone Write, and also scales to high write throughput under FEMU emulation [49]. When writes are issued to multiple zones per ZNS SSD, ZapRAID always maintains the highest throughput and lowest tail latency compared with the exclusive use of either Zone Append or Zone Write. Furthermore, ZapRAID maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 2.1 Zoned Namespace (ZNS)
The Zoned Namespace (ZNS) interface [14, 19] is a new NVMe Command Set [4] that exposes flash-based SSD internals to the host for host-level storage management. Figure 1 shows the architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model, which we use in this paper and also previous studies (e.g., [18, 42, 65]). The ZNS interface abstracts a ZNS SSD as append-only zones, each of which has a maximum size called the zone capacity. Each zone organizes data in logical blocks (or blocks in short in this paper), and each block is mapped to a physical flash page. Each flash page also has an out-of-band area for metadata storage (of size tens of bytes) located alongside the page content. Note that the out-of-band area is also available in general NVMe SSDs, and its space is not included in the total zone capacity. In this paper, we assume that the block size is 4 KiB and the out-of-band area size is 64 bytes, both of which can be configured in the ZN540 model when a ZNS SSD is formatted.
@@ 2.2 Zone Write versus Zone Append
Single-zone writes.
@@
Multi-zone writes.
@@
We observe that for small writes (e.g., of size 4 KiB or 8 KiB), Zone Append achieves higher write throughput than Zone Write in the single-zone setting by exploiting intra-zone parallelism, while Zone Write scales better and achieves higher write throughput than Zone Append when the number of open zones increases by better exploiting inter-zone parallelism. For large writes (e.g., of size 16 KiB), both Zone Write and Zone Append have similar write throughput for different numbers of open zones. A RAID system based on ZNS SSDs should adequately combine Zone Append and Zone Write to accommodate the scenarios with different numbers of open zones. Even though Zone Append only shows high performance when write requests are issued to a single zone, it is still beneficial for scenarios where shared storage systems need performance isolation [23, 32, 38]. Specifically, since a zone is typically mapped to only a subset of physical flash channels or chips [17], a shared storage system based on ZNS SSDs should allocate only a limited number of open zones to applications that do not have high write bandwidth demands. This isolates the write traffic of the applications to a subset of flash channels and limits their impact on other co-located applications. Thus, Zone Append can help such applications achieve high write performance through intra-zone parallelism and sustain the bursts of a large number of write requests if necessary, even with only a single zone per ZNS SSD being used.
Also, a RAID system based on ZNS SSDs should accommodate different request sizes. Workloads with mixed request sizes are common in practice; for example, practical cloud block storage workloads have 60% of write requests that are at most 4 KiB and 25% of write requests that are at least 16 KiB [51].
@@ 3.1 Design Overview
Each segment corresponds to $k+m$ zones in $k+m$ drives. In ZapRAID, each segment comprises three regions that span across $k+m$ drives: the header region and the footer region for keeping metadata for crash consistency (§3.4), and the data region for storing data and parity chunks. Specifically, the header region stores the zone IDs of all zones from the $k+m$ drives in the segment, the chunk size, the stripe group size (defined in §3.2), and the RAID scheme (e.g., RAID-5/RAID-6, $k$, and $m$). The footer region keeps the block metadata, including the LBA, write timestamp, and stripe ID of each block in the data region.
All three regions have pre-specified sizes. The header region contains exactly one stripe of $k+m$ chunks, each of which resides at the beginning of a zone. The data region contains a fixed number of stripes, denoted by $S$, following the header region. For each segment, each chunk contains a fixed number of blocks, denoted by $C$, so the number of blocks in the data region is $S\cdot C$. Suppose that the LBA size is 8 bytes, the write timestamp size is 8 bytes, the stripe ID size is 4 bytes, and the block size is 4 KiB. Thus, each block in the footer region can store the block metadata of $\lfloor\tfrac{4096}{20}\rfloor=204$ blocks, so the footer region occupies $\lceil\tfrac{S\cdot C}{204}\rceil$ blocks following the data region. For example, the size of a zone in a ZN540 ZNS SSD [13] is 1,077 MiB (or equivalently, 275,712 blocks). If the chunk size is 4 KiB (i.e., $C=1$), then the header, data, and footer regions occupy 1 block, 274,366 blocks, and 1,345 blocks in a zone, respectively.
@@
ZapRAID keeps a number of in-memory in-flight stripes for newly written blocks. Each in-flight stripe is kept in memory until all of its $k$ data chunks and $m$ parity chunks are formed and persisted. To maintain durability, ZapRAID acknowledges the writes of an in-flight stripe only after the whole in-flight stripe is persisted (note that acknowledging the write of each block can lead to data loss if a drive storing an acknowledged block fails but the parity chunks are yet generated).
@@ 3.3 Hybrid Data Management
To determine if a write request is issued to a small-chunk or large-chunk segment, ZapRAID examines whether the size of the write request is smaller than $C_{l}$. If so, ZapRAID assigns the write request to a small-chunk segment; otherwise, it assigns the write request to a large-chunk segment. Also, during garbage collection (§4), the rewritten data is written to a large-chunk segment to mitigate I/O overhead.
@@ 3.5 Complete Workflows
ZapRAID needs to handle two specific cases during writes. First, if the written stripe is the first stripe in a stripe group, ZapRAID should ensure that all the stripes in the previous stripe group are completely persisted; if the stripe group is also the first one of a segment, the header region of the segment should be completely persisted. Second, if the stripe group is also the last one in the data region, ZapRAID writes the block metadata of all blocks in the segment into the footer region. It also creates a new open segment and writes the segment information to the header region, so that the new open segment can serve new writes.
@@ 4 Implementation
We prototype ZapRAID as a user-space block device module in C++ with around 9.7 K LoC based on the Storage Performance Development Kit (SPDK) [8], which provides a set of C libraries for high-performance I/Os. Our prototype decomposes request handling into seven tasks, namely dispatch, device I/O, completion, indexing, encoding, segment state tracking, and cleaning.
Dispatch task.
@@
The dispatch task is responsible for processing write requests and updating the header and footer regions. For write requests, the dispatch task assigns the written blocks to segments and submits the I/O events for the written blocks to the device I/O task. If a block fills a stripe in a segment, the dispatch task also sends a parity generation event to the encoding task, which computes parity chunks and submits the I/O events for the parity chunks to the device I/O task. Recall that the writes of the first stripe of a stripe group need to wait for the previous stripe group to be completely persisted (§3.5); in this case, the dispatch task will suspend the request and retry later until the segment state is updated by the segment state tracking task (see below).
@@
For read requests, the dispatch task first sends an L2P table query event to the indexing task, which asynchronously notifies the dispatch task to issue the read requests upon retrieving PBAs for the requested blocks. The dispatch task then submits the I/O events to the device I/O task to retrieve the requested blocks from storage.
Device I/O task.
@@
The device I/O task is responsible for submitting block I/O requests to the drives. It receives I/O events from other tasks and submits the corresponding I/O requests to the drives. It is also responsible for checking the completion of each block I/O request from the drives and notifying the completion task (see below).
Completion task.
@@
The completion task tracks the completion state of each I/O request. It receives the notification of completion of a block I/O event from the device I/O task. If all the I/O events for a request are completed, the completion task notifies the upper-layer application. The application can register a callback function for each request. For write requests, the completion task first sends an L2P table update event to the indexing task, which triggers the callback function after updating the L2P table. For read requests, the completion task directly triggers the callback function.
@@
If a block is lost or unavailable during a read request, the completion task issues a degraded read to the block. It examines the compact stripe table to identify the PBAs of the available blocks of the same stripe. It then sends the I/O events to the device I/O task to retrieve the available blocks. After all available blocks are retrieved, the completion task decodes the lost block and marks the read request as completed.
Indexing task.
@@
The indexing task is responsible for the L2P table maintenance. It receives L2P table query events from the dispatch task and returns the PBAs of requested blocks for read requests. It also receives L2P table update events from the completion task and the cleaning task. If an L2P table update event is from the completion task, it acknowledges the write requests to the application on behalf of the completion task after the L2P table is updated. In addition, the indexing task is responsible for the offloading of L2P table entries if the L2P table size reaches the memory size limit (§3.1).
Encoding task.
@@
The encoding task is responsible for parity generation. When it receives a notification from the dispatch task that a stripe has $k$ data chunks, it generates parity chunks. It then submits I/O events for the parity chunks to the device I/O task. We handle parity chunk generation outside of the normal I/O requests to mitigate interference on normal I/O requests.
Segment state tracking task.
@@
The segment state tracking task examines whether the system needs to write the header region or the footer region of a segment, or whether it has persisted all stripes in a stripe group (§3.5). It periodically (every 1 $\mu$s in our prototype) examines the segment states. When the segment state needs to be updated, it also sends the corresponding I/O events to the device I/O task.
Cleaning task.
@@
The cleaning task is responsible for reclaiming the space of stale blocks. It periodically (every 1 $\mu$s in our prototype) examines the available space and triggers garbage collection if the available space drops below a threshold. It selects one sealed segment that has the most stale blocks and rewrites its valid blocks by submitting block I/O events to the device I/O task. Currently, our prototype tracks the validity of all blocks (including both user-written blocks and mapping blocks) for each segment via an in-memory bitmap (whose content can also be derived from both the L2P table for user-written blocks and the LBA-to-PBA table for mapping blocks) and collects the LBAs of valid blocks from the block metadata. When all rewrites are completed, the dispatch task sends the LBAs of all rewritten blocks, along with their old and new PBAs, to the indexing task for updating the L2P table.
@@
We assign the seven tasks to SPDK threads, which exchange events via spdk_thread_send_msg and event polling. Currently, we co-locate the dispatch task, completion task, segment state tracking task, and cleaning task in the main SPDK thread. We also assign the device I/O task, indexing task, and encoding task separately to three different SPDK threads for parallelism, such that our prototype can saturate the write bandwidth of our testbed.

### v2

@@ Abstract
Zoned Namespace (ZNS) defines a new abstraction for host software to flexibly manage storage in flash-based SSDs as append-only zones. It also provides a Zone Append primitive to further boost the write performance of ZNS SSDs by exploiting intra-zone parallelism. However, making Zone Append effective for reliable and scalable storage, in the form of a RAID array of multiple ZNS SSDs, is non-trivial since Zone Append offloads address management to ZNS SSDs and requires hosts to dedicatedly manage RAID stripes across multiple drives. We propose ZapRAID, a high-performance log-structured RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write parallelism and lightweight stripe management. ZapRAID adopts a group-based data layout with a coarse-grained ordering across multiple groups of stripes, such that it can use small-size metadata for stripe management on a per-group basis under Zone Append. It further adopts hybrid data management to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Write and Zone Append primitives. We implement ZapRAID as a user-space block device, and evaluate ZapRAID using microbenchmarks, trace-driven experiments, and real-application experiments. Our evaluation results show that ZapRAID achieves high write throughput and maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 1 Introduction
ZNS SSDs support two write primitives, namely Zone Write and Zone Append, where Zone Append provides opportunities for further performance gains by exploiting intra-zone parallelism. Specifically, a ZNS SSD tracks per-zone write pointers, such that any write to a zone must specify the same offset indicated by the write pointer. Zone Write requires the host to specify the block address when writing a block as in conventional SSDs. This provides backward compatibility with existing applications. However, in order to match the offset of a Zone Write with the on-device write pointer of a zone, the host can only issue one request to each zone at a time, thereby limiting intra-zone parallelism. In contrast, Zone Append eliminates the need of the host to specify block addresses in writes by fully offloading address management to ZNS SSDs. In Zone Append, the host only specifies the zone to which a write is issued, and the ZNS SSD returns the address to the host upon write completion. Thus, the host can issue multiple Zone Append commands within one zone to exploit the intra-zone parallelism for improved write performance.
Despite the performance gains of Zone Append, its offloading of address management to ZNS SSDs implies that the host not only cannot directly specify the block addresses of writes in a zone, but it also cannot control the ordering of writes in concurrent Zone Append commands. This creates new challenges to storage system designs that require host-level address management, particularly when applying Redundant Array of Independent Disks (RAID) [64] to form an array of multiple ZNS SSDs for reliable and scalable storage. Specifically, in a traditional RAID array, a RAID controller runs atop multiple drives and organizes data in stripes, each of which encodes a set of data blocks into parity blocks and distributes the data and parity blocks across drives. For efficient data repair, the RAID controller statically assigns the same block address in each drive for the blocks of the same stripe, so that the repair of any lost block can directly retrieve the available data and parity blocks of the same stripe from other drives for decoding. However, under Zone Append, it is infeasible for the RAID controller to manage block addresses in a zone; instead, the RAID controller needs to maintain dedicated address mapping information to specify the block locations for each stripe, which unavoidably incurs performance penalties to stripe management.
@@
From our measurement study (§2.2), we observe that Zone Append effectively exploits intra-zone parallelism when a workload is dominated by small writes and when writes are issued to a single zone of a ZNS SSD at a time, while Zone Write is beneficial when a workload is dominated by large writes or when writes are concurrently issued to multiple zones of a ZNS SSD (in which Zone Write exploits better inter-zone parallelism for large writes). To this end, ZapRAID adopts hybrid data management to manage small and large writes across multiple zones in a ZNS SSD, so as to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Append and Zone Write primitives.
@@
  • We design ZapRAID, a high-performance Log-RAID system for ZNS SSDs. ZapRAID builds on two core ideas: (i) a group-based data layout under Zone Append (§3.2) and (ii) hybrid data management (§3.3) under a combination of Zone Write and Zone Append. ZapRAID further addresses reliability in terms of metadata persistence and crash consistency (§3.4).
  • We prototype ZapRAID as a user-space block device with the Storage Performance Development Kit (SPDK) [11], such that ZapRAID exports a block interface for compatibility with general applications and can parallelize its internal operations with multiple SPDK threads.
  • We evaluate our ZapRAID prototype on real ZNS SSD devices using microbenchmarks, trace-driven experiments, and real-application experiments. When writes are issued to a single zone per ZNS SSD, ZapRAID increases the write throughput by up to 77.2% and reduces the tail latency by up to 36.4% compared with the exclusive use of Zone Write, and also scales to high write throughput under FEMU emulation [53]. When writes are issued to multiple zones per ZNS SSD, ZapRAID always maintains the highest throughput and lowest tail latency compared with the exclusive use of either Zone Write or Zone Append. Furthermore, ZapRAID maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 2.1 Zoned Namespace (ZNS)
The Zoned Namespace (ZNS) interface [15, 21] is a new NVMe Command Set [4] that exposes flash-based SSD internals to the host for host-level storage management. Figure 1 shows the architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model, which we use in this paper and also previous studies (e.g., [20, 45, 49, 69, 32]). The ZNS interface abstracts a ZNS SSD as append-only zones, each of which has a maximum size called the zone capacity. Each zone organizes data in logical blocks (or blocks in short in this paper), and each block is mapped to a physical flash page. Each flash page also has an out-of-band area for metadata storage (of size tens of bytes) located alongside the page content. Note that the out-of-band area is also available in general NVMe SSDs, and its space is not included in the total zone capacity. In this paper, we assume that the block size is 4 KiB and the out-of-band area size is 64 bytes, both of which can be configured in the ZN540 model when a ZNS SSD is formatted.
@@ 2.2 Zone Write versus Zone Append
Writes to a single open zone.
@@
Writes to multiple open zones.
@@
We observe that for small writes (e.g., of size 4 KiB or 8 KiB), Zone Append achieves higher write throughput than Zone Write in the single-zone setting by exploiting intra-zone parallelism, while Zone Write scales better and achieves higher write throughput than Zone Append when the number of open zones increases by better exploiting inter-zone parallelism. For large writes (e.g., of size 16 KiB), both Zone Write and Zone Append have similar write throughput for different numbers of open zones. A RAID system based on ZNS SSDs should adequately combine Zone Append and Zone Write to accommodate the scenarios with different numbers of open zones. Even though Zone Append only shows high performance when write requests are issued to one or a few open zones, it is still beneficial for the following scenarios. First, shared storage systems need performance isolation [25, 35, 41]. Specifically, since a zone is typically mapped to only a subset of physical flash channels or chips [19], a shared storage system based on ZNS SSDs should allocate only a limited number of open zones to applications that do not have high write bandwidth demands. This isolates the write traffic of the applications to a subset of flash channels and limits their impact on other co-located applications. Second, it is common to have mixed request sizes in many real-world workloads. For example, practical cloud block storage workloads have 60% of write requests that are at most 4 KiB and 25% of write requests that are at least 16 KiB [55]. By separating small and large writes into different open zones (§3.3), we can use Zone Append and Zone Write for different groups of writes to improve the overall performance (§5.2). Thus, Zone Append can help such applications achieve high write performance through intra-zone parallelism and sustain the bursts of a large number of write requests if necessary, even with only one or a few open zones per ZNS SSD being used.
@@ 3.1 Design Overview
Each segment corresponds to $k+m$ zones in $k+m$ drives. In ZapRAID, each segment comprises three regions that span across $k+m$ drives: the header region and the footer region for keeping metadata for crash consistency (§3.4), and the data region for storing data and parity chunks. Specifically, the header region stores the RAID scheme (including RAID-5/RAID-6, $k$, and $m$) and the segment-specific information (including the zone IDs of all zones from the $k+m$ drives in the segment, the chunk size, and the stripe group size (defined in §3.2)). The footer region keeps the block metadata, including the LBA, write timestamp, and stripe ID of each block in the data region.
All three regions have pre-specified sizes. The header region contains exactly one stripe of $k+m$ chunks, each of which resides at the beginning of a zone. The data region contains a fixed number of stripes, denoted by $S$, following the header region. For each segment, each chunk contains a fixed number of blocks, denoted by $C$, so the number of blocks in the data region is $S\cdot C$. Suppose that the LBA size is 8 bytes, the write timestamp size is 8 bytes, the stripe ID size is 4 bytes, and the block size is 4 KiB. Thus, each block in the footer region can store the block metadata of $\lfloor\tfrac{4096}{20}\rfloor=204$ blocks, so the footer region occupies $\lceil\tfrac{S\cdot C}{204}\rceil$ blocks following the data region. For example, the capacity of a zone in a ZN540 ZNS SSD [14] is 1,077 MiB (or equivalently, 275,712 blocks). If the chunk size is 4 KiB (i.e., $C=1$), then the header, data, and footer regions occupy 1 block, 274,366 blocks, and 1,345 blocks in a zone, respectively.
@@
ZapRAID keeps a number of in-memory in-flight stripes for newly written blocks. Each in-flight stripe is kept in memory until all of its $k$ data chunks and $m$ parity chunks are formed and persisted. To maintain durability, ZapRAID acknowledges the writes of an in-flight stripe only after the whole in-flight stripe is persisted (note that acknowledging the write of each block can lead to data loss if a drive storing an acknowledged block fails but the parity chunks are not yet generated).
@@ 3.3 Hybrid Data Management
To determine if a write request is issued to a small-chunk or large-chunk segment, ZapRAID examines whether the size of the write request is smaller than a write request size threshold, which we now set as $C_{l}$. If so, ZapRAID assigns the write request to a small-chunk segment; otherwise, it assigns the write request to a large-chunk segment. Also, during garbage collection (§4), the rewritten data is written to a large-chunk segment to mitigate I/O overhead.
@@ 3.5 Complete Workflows
ZapRAID needs to handle two specific cases during writes. First, if the written stripe is the first stripe in a stripe group, ZapRAID should ensure that all the stripes in the previous stripe group are completely persisted; if the stripe group is also the first one of a segment, the header region of the segment should be completely persisted. Second, if the stripe group is also the last one in the data region, ZapRAID writes the block metadata of all blocks in the segment into the footer region in the background. It also creates a new open segment and writes the segment information to the header region (without waiting for the completion of the writing of the footer region in the previous segment), so that the new open segment can serve new writes.
@@
3.6 Discussion
We now discuss some open issues of our ZapRAID design.
Choice of the write request size threshold in hybrid data management.
Recall that ZapRAID compares the size of a write request against a threshold to determine if the write request is assigned to a small-chunk or large-chunk open segment (§3.3). Currently, the threshold is set to $C_{l}$. The choice of the threshold poses different performance trade-offs. A larger threshold implies that more writes will go to small-chunk segments. Full small-chunk stripes are formed faster, and hence small writes are also acknowledged faster (i.e., lower response time). However, the write throughput is lower since writing small-chunk segments is slower than writing large-chunk segments (based on the results in Figure 2 for different write request sizes). For simplicity, we assume that the threshold is equal to $C_{l}$ (i.e., if a write request size is smaller than $C_{l}$, it is assigned to a small-chunk segment; otherwise, it is assigned to a large-chunk segment).
Relationships between the ZapRAID design and the ZNS SSD model.
While our motivating experiments and evaluation are based on the ZN540 model [14], our ZapRAID design does not heavily depend on a specific ZNS SSD model. For example, some of our key design features, such as the group-based data layout and crash recovery management, apply to all ZNS SSD models.
ZapRAID is mainly based on two design principles: (i) Zone Append outperforms Zone Write in one or a few open zones for small writes, and (ii) a combination of Zone Append and Zone Write can be used for hybrid data management when there are multiple open zones. We note that ZapRAID now only uses Zone Append for a single open segment in hybrid data management since the improvement of Zone Append is limited for a large number of open zones, and we conjecture that the reason is due to the computational overhead of the firmware. Nevertheless, if the computational overhead can be mitigated in future ZNS SSD models, ZapRAID can issue Zone Append to more open segments to better exploit intra-zone parallelism; this can be readily achieved with parameter reconfigurations based on our current design.
Our ZapRAID design is also applicable to different zone capacities of ZNS SSDs. For smaller zone capacities, the data and footer regions become smaller, yet the group-based data layout remains applicable. For example, when the zone capacity is 96 MiB [62] (or equivalently, 24,576 blocks), for 4-KiB chunks, the header, data, and footer regions occupy 1 block, 24,455 blocks, and 120 blocks, respectively (based on the calculation of §3.1). If the stripe group size is 256, there are still many stripe groups (i.e., 96 groups) for intra-zone parallelism.
Note that the ZN540 model has been widely studied by a large body of prior work in the literature (e.g., [20, 45, 49, 69, 32]), and hence is representative for ZNS SSD research.
@@ 4 Implementation
Since ZNS SSDs offer low latency and high throughput, the overhead of the Linux kernel I/O stack (e.g., high latencies of interrupts, locks, and system calls) is no longer negligible [50]. To reduce I/O latency, we implement ZapRAID using the Storage Performance Development Kit (SPDK) [9]. SPDK supports polling-based and lockless I/Os. It also avoids system calls by moving I/O drivers to the user space.
@@
Specifically, we prototype ZapRAID as a user-space block device module in C++ with around 9.7 K LoC based on SPDK. ZapRAID exports a block interface to be compatible with general applications. Our prototype decomposes request handling into seven handlers, namely dispatch, device I/O, completion, indexing, encoding, segment state tracking, and cleaning.
Dispatch handler.
@@
The dispatch handler is responsible for processing write requests and updating the header and footer regions. For write requests, the dispatch handler assigns the written blocks to segments and submits the I/O events for the written blocks to the device I/O handler. If a block fills a stripe in a segment, the dispatch handler also sends a parity generation event to the encoding handler, which computes parity chunks and submits the I/O events for the parity chunks to the device I/O handler. Recall that the writes of the first stripe of a stripe group need to wait for the previous stripe group to be completely persisted (§3.5); in this case, the dispatch handler will suspend the request and retry later until the segment state is updated by the segment state tracking handler (see below).
@@
For read requests, the dispatch handler first sends an L2P table query event to the indexing handler, which asynchronously notifies the dispatch handler to issue the read requests upon retrieving PBAs for the requested blocks. The dispatch handler then submits the I/O events to the device I/O handler to retrieve the requested blocks from storage.
Device I/O handler.
@@
The device I/O handler is responsible for submitting block I/O requests to the drives. It receives I/O events from other handlers and submits the corresponding I/O requests to the drives. It is also responsible for checking the completion of each block I/O request from the drives and notifying the completion handler (see below).
Completion handler.
@@
The completion handler tracks the completion state of each I/O request. It receives the notification of completion of a block I/O event from the device I/O handler. If all the I/O events for a request are completed, the completion handler notifies the upper-layer application. The application can register a callback function for each request. For write requests, the completion handler first sends an L2P table update event to the indexing handler, which triggers the callback function after updating the L2P table. For read requests, the completion handler directly triggers the callback function.
@@
If a block is lost or unavailable during a read request, the completion handler issues a degraded read to the block. It examines the compact stripe table to identify the PBAs of the available blocks of the same stripe. It then sends the I/O events to the device I/O handler to retrieve the available blocks. After all available blocks are retrieved, the completion handler decodes the lost block and marks the read request as completed.
Indexing handler.
@@
The indexing handler is responsible for the L2P table maintenance. It receives L2P table query events from the dispatch handler and returns the PBAs of requested blocks for read requests. It also receives L2P table update events from the completion handler and the cleaning handler. If an L2P table update event is from the completion handler, it acknowledges the write requests to the application on behalf of the completion handler after the L2P table is updated. In addition, the indexing handler is responsible for the offloading of L2P table entries if the L2P table size reaches the memory size limit (§3.1).
Encoding handler.
@@
The encoding handler is responsible for parity generation. When it receives a notification from the dispatch handler that a stripe has $k$ data chunks, it generates parity chunks. It then submits I/O events for the parity chunks to the device I/O handler. We handle parity chunk generation outside of the normal I/O requests to mitigate interference on normal I/O requests.
Segment state tracking handler.
@@
The segment state tracking handler examines whether the system needs to write the header region or the footer region of a segment, or whether it has persisted all stripes in a stripe group (§3.5). It periodically (every 1 $\mu$s in our prototype) examines the segment states. When the segment state needs to be updated, it also sends the corresponding I/O events to the device I/O handler.
Cleaning handler.
@@
The cleaning handler is responsible for reclaiming the space of stale blocks. It periodically (every 1 $\mu$s in our prototype) examines the available space and triggers garbage collection if the available space drops below a fixed threshold. It selects one sealed segment that has the most stale blocks and rewrites its valid blocks by submitting block I/O events to the device I/O handler. Currently, our prototype tracks the validity of all blocks (including both user-written blocks and mapping blocks) for each segment via an in-memory bitmap (whose content can also be derived from both the L2P table for user-written blocks and the LBA-to-PBA table for mapping blocks) and collects the LBAs of valid blocks from the block metadata. When all rewrites are completed, the dispatch handler sends the LBAs of all rewritten blocks, along with their old and new PBAs, to the indexing handler for updating the L2P table.
In this work, we do not further consider sophisticated garbage collection algorithms, which need to address the key design issues of (i) when to trigger garbage collection, (ii) which segment to clean for garbage collection, and (iii) where to rewrite the valid data. We argue that such design issues are orthogonal to our current focus on leveraging Zone Append and Zone Write for a high-performance RAID design. We pose the design of new garbage collection algorithms as a future work.
@@
We assign the seven handlers to SPDK threads, which exchange events via the function spdk_thread_send_msg and event polling. Currently, we co-locate the dispatch handler, completion handler, segment state tracking handler, and cleaning handler in the main SPDK thread. We also assign the device I/O handler, indexing handler, and encoding handler separately to three different SPDK threads for parallelism, such that our prototype can saturate the write bandwidth of our testbed.